The graph of y = x^4 + 7x^2 - 8 is symmetric with respect to the y axis.
a. True
b. False

Answer by stanbon(75887)

The graph of y = x^4 + 7x^2 - 8 is symmetric with respect to the y axis.
a. True
b. False
--------------
y-symmetry????
f(x)=x^4+7x^2-8
f(-x)=(-x)^4+7(-x)^2-8= x^4+7x^2-8
Since f(x)=f(-x) there is y-axis symmetry.
---------------------------------
Cheers,
stan H.